OpenStudy (anonymous):

E? Does the river speed play a factor? David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing?

OpenStudy (anonymous):

y = x + 4 y = x − 4 y = 6 − 2x y = 6 − x y = 4 − x

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

I thought so

OpenStudy (igreen):

No, I think that's wrong..plus you're not supposed to give answers.

OpenStudy (igreen):

6 - 2 = ?

OpenStudy (anonymous):

4, how do I get you to be a QH you're good at everything haah

OpenStudy (igreen):

Okay, so he's traveling at about 4 MPH since he's going 6 and the river is against him at 2 miles per hour. Then it says "as he gets tired his speed decreases by 1 MPH. So we subtract 1x to it, which is the same as 'x'.

OpenStudy (anonymous):

y=4-x

OpenStudy (igreen):

Yep, that's our answer.